b) A 20kg mass of Aluminum of specific heat Capacity OF 900kg /k was heated from 2k to 12k. Determine the


amount of heat Energy required


1
Expert's answer
2022-03-17T15:16:33-0400

"Q=cm\Delta{T}=900\frac{J}{kg\times{K}}\times20\ kg\times(12\ K-2\ K)=180000\ J=180\ kJ"


Answer: 180 kJ